sunny has some 10 rupee coins and some 5 rupee coins. He has 3 five rupee coins more than the number of 10 rupee coins. if he has a total of Rs.135, how many coins of each variety has he got
Answer by mananth(16946)

10 rupee coins be x
5 rupee coins will be x+3
Total value = Rs135
10x +5(x+3) =135
10x+5x+15 =135
20x= 135-15
20x=120
x= 120/12
x= 10
Ten rupee coins 10
5 rupee coins 10+3 = 13